### v1

# Security flaws from time-varying active encoding in high-speed measurement-device-independent quantum key distribution

Abstract
Quantum key distribution (QKD) can transmit secret keys with, in principle, information-theoretic security. However, bandwidth limitations in practical equipment threaten the security of high-speed (GHz) QKD systems. We propose and characterize a new side channel which arises when using active encoding. As an illustrative example, we focus on electro-optic phase modulation for polarization encoding at 1 GHz. We show that this side channel may reduce the maximum secure transmission distance by over 50% in a decoy state measurement-device-independent QKD protocol.
preprint: APS/123-QED
Introduction.— Quantum key distribution is a method which allows two parties to share a secret key with information-theoretic, quantum-safe security Xu et al. (2020). This is commonly done through encoding information into a train of quantum optical states. Unfortunately, practical implementations of QKD often fail to strictly uphold the security requirements which are posed by QKD theory for information-theoretic security. Imperfections come in the form of side channels which may leak information to an eavesdropper as well as state preparation flaws (SPFs) which are errors in the prepared quantum states Xu et al. (2020). As evidenced by various quantum hacking experiments, side channels present significant security risks to QKD protocols, making their characterization and mitigation vital areas of research Xu et al. (2020, 2010); Gottesman et al. (2004); Tang et al. (2013); Zhao et al. (2008); Lydersen et al. (2010); Brassard (2005); Jain et al. (2014). The development of the measurement-device-independent QKD (MDI QKD) protocol marked a significant milestone in QKD research by eliminating all side channels at the detector Lo et al. (2012). However, source side channels remain a major issue requiring further research Lo et al. (2012); Xu et al. (2020). Source side channels and imperfections could significantly lower the secret key rate in QKD Currás-Lorenzo et al. (2023); Tamaki et al. (2014) and therefore deserve our close attention.
The secret key rate per second is crucial for assessing QKD system performance. Rapid key transmission is practically useful, especially when using QKD alongside the one-time pad encryption technique, which requires the key length to match the message length for information-theoretic security Shannon (1949). As the secret key rate per pulse is fundamentally constrained by limits such as the PLOB bound Pirandola et al. (2017), enhancing the optical pulse transmission speed becomes pivotal for elevating the secret key rate. This has motivated a push towards high speed (GHz clock rate) QKD Roger et al. (2023); Boaron et al. (2018); Li et al. (2019); Grünenfelder et al. (2020); ichiro Yoshino et al. (2018); Woodward et al. (2021); Beutel et al. (2021); Chen et al. (2023); Lu et al. (2021). However, at high speeds, limited equipment bandwidth leads to the emergence of novel side channels at the source. The pattern effect is a notable example which has garnered significant interest ichiro Yoshino et al. (2018); Lu et al. (2021); Kang et al. (2023); Roberts et al. (2018); Pereira et al. (2020).
Here, we propose and characterize a new side channel which may pose a significant threat to the security of high speed QKD systems - time-varying encoding. Consider the standard BB84 QKD protocol, which requires each prepared state to be randomly chosen from one of four specific qubit states which are required for information-theoretic security Xu et al. (2020). At high clock rates, using bandwidth-limited modulation equipment to encode optical states can result in encoding variations across the state duration/width. Therefore, the optical states possess a temporal degree of freedom tied to the encoding process, which serves as a side channel. We emphasize that this would be an issue in any discrete-variable QKD protocol using bandwidth-limited active modulation for encoding.
This has motivated our theoretical and experimental study of this newly identified side channel. As an illustrative example, we focus on active BB84 polarization encoding, which is often achieved using an electro-optic phase modulator. We present simulated and experimental time-varying polarization encoding profiles corresponding to a repetition rate of 1 GHz. Using a numerical security proof based on semidefinite programming Primaatmaja et al. (2019), we use our simulated and experimental results to obtain the asymptotic key rate for a decoy state MDI QKD protocol. Our results demonstrate a significant (over 50%) reduction in the maximum secure transmission distance due to the time-varying encoding side channel.
Quantum Optical Modelling.— Here, we derive an expression for the single photon component of a quantum optical coherent state with time-varying polarization encoding. This will be used to compute the single photon phase error rate (used to bound an eavesdropper’s information on the secret key in a decoy state protocol Lo et al. (2005)). We will use |$\ket{\psi^{i}_{x}}$ to represent this single photon component, where $i$ represents the basis choice and $x$ represents the bit choice.
We start by expressing the time-varying polarization state of |$\ket{\psi^{i}_{x}}$. Note that active polarization encoding is typically achieved using an electro-optic phase modulator to modulate the phase along a specific polarization component. Without loss of generality, we consider an input (unmodulated) polarization state of |$\frac{\ket{H}+\ket{V}}{\sqrt{2}}$ (where |$\ket{H}$ and |$\ket{V}$ are orthogonal, linear polarization states). Now, let $\phi^{i}_{x}(t)$ refer to the time-varying phase change applied by the phase modulator to the |$\ket{V}$ component. The output time-varying polarization state would then be |$\frac{\ket{H}+e^{i\phi^{i}_{x}(t)}\ket{V}}{\sqrt{2}}$.
Now, we consider the optical intensity profile of the quantum optical coherent state. We assume a Gaussian intensity profile. Hence, the temporal probability distribution of the single photon component would follow a Gaussian function, $f(t)=ae^{-\frac{t^{2}}{2c^{2}}}$. We choose $a$ such that $\int_{-\infty}^{\infty}f(t)=1$. The parameter $c$ is related to the full width at half maximum (FWHM) of the optical pulse (recall that the FWHM of a Gaussian pulse is $2\sqrt{2\ln 2}c$).
Hence, |$\ket{\psi^{i}_{x}}$ can be written as follows (leaving out the global phase term, which we assume in this work to be time-independent).
[…]
Time-Varying Encoding is not Loss-Tolerant.— We can intuitively understand the level of information leakage inherent in |$\ket{\psi^{i}_{x}}$ using the formulation described in Currás-Lorenzo et al. (2023). We start by breaking down the state as follows.
|$\ket{\psi^{i}_{x}}=\sqrt{1-\epsilon^{i}_{x}}\ket{\theta^{i}_{x}}+\sqrt{\epsilon^{i}_{x}}\ket{\theta^{i\perp}_{x}}$ (2)
Here, |$\ket{\theta^{i}_{x}}$ is the closest (in terms of fidelity) qubit state to |$\ket{\psi^{i}_{x}}$. Here, we assume the form of such a state to match Equation 1 with one key difference - the polarization is not time-dependent (which results in separability between the temporal degree of freedom and polarization degree of freedom). Hence, we can express |$\ket{\theta^{i}_{x}}$ as follows, where $\theta^{i}_{x}$ represents a time-independent phase change applied by the phase modulator ($\theta^{i}_{x}$ is chosen such that $|\braket{\theta^{i}_{x}}{\psi^{i}_{x}}|^{2}$ is maximal).
[…]
The remaining component, |$\ket{\theta^{i\perp}_{x}}$ effectively represents the side channel within the transmitted state and is orthogonal to |$\ket{\theta^{i}_{x}}$. The parameter $\epsilon^{i}_{x}$ represents the level of deviation of |$\ket{\psi^{i}_{x}}$ from a qubit state due to time-varying encoding. It can be expressed as follows.
$|\braket{\theta^{i}_{x}}{\psi^{i}_{x}}|^{2}=1-\epsilon^{i}_{x}$ (4)
Note that the imperfection introduced by time-varying encoding differs from a simple SPF, which would only introduce a polarization error while preserving the encoded state as a qubit. As demonstrated by the loss-tolerant protocol for bounding secret key rates in QKD, SPFs do not render a QKD protocol vulnerable to channel loss Tamaki et al. (2014); Pereira et al. (2019). However, this is not the case for the side channel introduced by a non-qubit component of the encoded state Currás-Lorenzo et al. (2023). Eve could potentially exploit channel loss to extract information about the encoded states through this side channel (by performing an unambiguous state discrimination attack, for example), posing significant threats to the secret key rate and maximum secure transmission distance Currás-Lorenzo et al. (2023). The impact of the side channel on security can be understood through the parameter $\epsilon^{i}_{x}$. We will later use this parameter in order to further understand and contextualize our secret key rate results.
Characterization of Time-Varying Phase Modulation.— Here, we acquire $\phi^{i}_{x}(t)$ (which we will refer to as “temporal phase profiles” or simply “phase profiles”) in order to concretely understand the level of time-variance in high-speed bandwidth-limited electro-optic phase modulation when using our state-of-the-art lab equipment.
We start by reviewing electro-optic phase modulation Gee et al. (1983). An electro-optic phase modulator works through a traveling wave mechanism in which an optical signal and traveling RF signal propagate together through the modulator with their speeds matched. The resulting temporal phase profile of the optical signal is proportional to the temporal voltage profile of the RF signal. The voltage required to create a $\pi$ phase change is referred to as $V_{\pi}$ (more generally, we will use $V_{\phi}$ to refer to the voltage required to create a $\phi$ phase change).
Recall that in this paper, we consider BB84 encoding. Hence, we are interested in the temporal phase profiles resulting from the following phase modulations - $\pi/2$, $\pi$, and $-\pi/2$. The width of the phase modulation window (i.e. the width of the voltage pulse used to modulate phase) must be adequately larger than the width of the optical pulse being modulated. Here, we consider a 50 ps wide optical pulse and a 200 ps wide phase modulation window, typically used at GHz clock rates Lu et al. (2021); Kang et al. (2023); Roberts et al. (2018); ichiro Yoshino et al. (2018); Li et al. (2019). Ideally, the voltage pulse used to modulate phase should be square (or have a flat modulation region at the peak), resulting in time-invariant modulation. However, the bandwidth limitations of the arbitrary waveform generator (AWG) and RF amplifier (which is often necessary to create voltage pulses with the required amplitude) result in distorted voltage pulses, deviating from a perfect square shape.
While limited bandwidth is a notable imperfection impacting the time-variance of phase modulation, other equipment-specific factors, such as nonlinearity and the frequency response function, may also contribute. This motivates us to provide optimistically simulated phase profiles in addition to our measured phase profiles as a generalized baseline. The simulated phase profiles solely consider the bandwidth limitations of the AWG, RF amplifier, and electro-optic modulator. We optimistically approximate the frequency response of each component using a simple Butterworth filter of order 2, with -3 dB points of 25 GHz, 11 GHz, and 15 GHz, respectively (corresponding to the equipment specifications). Butterworth filters are an optimistic choice, as their frequency response in the passband is relatively flat. Furthermore, at an order of 2, the frequency response falls off relatively gradually past the passband (in comparison to the frequency response of our state-of-the-art equipment) key (2017); drv (2023); jds (2004). These filters were created and applied to a 200 ps wide square pulse in order to generate the simulated profile (using the SciPy signal processing class) Contributors (2022). The result can be seen in Figure 1.
Figure 1: Simulated $\pi$ phase profile (the simulated ±$\pm\pi/2$ phase profiles are simply a scaled version of this). This was generated by applying Butterworth filters to a 200 ps wide square pulse. The Butterworth filters represent the AWG, amplifier, and electro-optic modulator. The optical pulse intensity profile assumed in this work (50 ps FWHM Gaussian pulse) is also shown in green.
We now move to our experimentally acquired temporal phase profiles. The experimental setup for generating and measuring the phase profiles is shown in Figure 2. The 25 GHz AWG was set to generate 200 ps wide (FWHM) square voltage pulses (the pulse repetition frequency was set to 10 MHz to reduce pulse correlations down to negligible levels). This pulse train was then amplified by an 11 GHz amplifier so that the voltage pulses have the necessary amplitude to generate the desired phase change. Using this setup, voltage pulses with an amplitude of $V_{\pi}$, $V_{\pi/2}$, and $V_{-\pi/2}$ were generated in order to acquire the $\pi$, $\pi/2$, and $-\pi/2$ temporal phase profiles. For our electro-optic modulator, $V_{\pi}$, $V_{\pi/2}$, and $V_{-\pi/2}$ are approximately 5.5 V, 2.75 V, and -2.75 V, respectively.
Figure 2: Experimental setup for generating and measuring phase profiles. The 25 GHz Keysight M8195A AWG (AWG) and 11 GHz ixBlue DR-VE-10-MO RF amplifier are used to generated voltage pulses which are sent to the JDSU OC-192 intensity modulator (IM). This intensity modulator acts as a phase modulator with a built in Mach-Zehnder interferometer to convert phase changes to intensity changes. The optical intensity profile (related to the optical phase profile through Equation 5) is collected using the Newport D-8ir 8 ps photodiode (PD) and the 12 GHz Agilent DSO81204A oscilloscope (OSC) after being amplified by a Thorlabs semiconductor optical amplifier (OA). The OA and PD were operated within their linear response region. Yellow indicates an optical fiber connection.
The $\pi$, $\pi/2$, and $-\pi/2$ temporal phase profiles were each generated by using the corresponding amplified voltage pulses to drive an electro-optic intensity modulator. An electro-optic intensity modulator uses electro-optic phase modulation in conjunction with Mach-Zehnder interferometry in order to modulate intensity. The phase change ($\phi$) applied by the intensity modulator is related to the normalized output intensity ($I$) through Equation 5 when the intensity modulator is biased at a minimum. This relationship can be used to determine a phase profile from its corresponding intensity profile.
$\phi=2cos^{-1}(\sqrt{1-I})$ (5)
Figure 3: a),b) Voltage pulses used for polarization/phase modulation as well as the corresponding phase profiles. Cubic spline interpolations of the profiles are shown using a solid line. The optical pulse intensity profile assumed in this work (50 ps FWHM Gaussian pulse) is also shown in green. Each profile was measured 10 times over an extended time period of several minutes (each time, oscilloscope averaging was used and set to 1024 to average out random noise, such as electronic noise) - the error bars represent the standard deviation in the 10 trials. c) Scaled version of a) demonstrating the difference between the voltage profiles used to modulate the two states. d) Scaled version of b) demonstrating the difference in temporal profiles between the two states.
The resulting temporal phase profiles are shown in Figure 3 (excluding the $-\pi$/2 profile, which we found to be nearly identical to the $\pi/2$ profile in shape). The corresponding voltage profiles are also shown. Evidently, the $\pi$ and $\pi/2$ phase profiles have a significant level of temporal variation (especially in comparison to the optimistically simulated phase profile from Figure 1, as expected). Interestingly, they do not have the same shape. We see the same effect in the corresponding voltage profiles. Likely, this results from the nonlinear (with respect to input voltage) response of the amplifier near its 1 dB compression point com of 5.6 V drv (2023).
In the next section, we will use our experimentally measured phase profiles in order to demonstrate the significant impact which time-varying polarization encoding has on the secret key rate in a decoy state MDI QKD protocol. We will also use our optimistically simulated phase profile to illustrate the extent of the impact, focusing solely on bandwidth limitations and excluding other practical considerations such as nonlinearity and imperfect frequency response.
Key Rate Results.— A symmetric polarization encoding (BB84) decoy state MDI QKD protocol with intensity settings of 0.05, 0.1, and 0.6 (near typical values Ma et al. (2012)) was assumed in the key rate analysis. Without loss of generality, we consider a projection onto the |$\ket{\Psi^{+}}$ Bell state as a passing measurement. A fiber loss of 0.2 dB/km and a dark count rate of $10^{-6}$/pulse was assumed. Moreover, the following simplifying assumptions were made to isolate specifically for the impact of time-varying encoding on the secret key rate. Although we neglect certain imperfections, they can be considered in a straightforward manner (unless stated otherwise) by deriving the observables used in the key rate formula accordingly and/or modifying the quantum optical model for the transmitted states.
  • Perfect polarization alignment to the electro-optic phase modulator (halfway between the fast and slow axes). Realistically, a certain degree of misalignment would be present.
  • A Gaussian optical pulse profile.
  • No temporal variation of global phase within the phase randomization window.
  • The temporal phase profiles of Bob’s states are the same as the temporal phase profiles of Alice’s states. In reality, Bob’s equipment specifications may be different from Alice’s equipment specifications. This may result in Bob’s states having different temporal phase profiles.
In our previous work Bourassa et al. (2022), we demonstrate the usage of a tight numerical key rate analysis technique based on semidefinite programming Primaatmaja et al. (2019) to deal with a specific time-dependent side channel. We show that the unique ability of the proof technique to deal with the exact time-dependent nature of the transmitted states (rather than a simplified, time-independent pessimistic approximation) results in a significantly tighter bound on the secret key rate. Hence, we chose to employ this same technique to analyze the impact of time-varying encoding on the secret key rate. Details regarding the technique can be found in the Supplementary Material.
At a high level, this technique uses the inner products between the transmitted states (i.e the Gram matrix of the transmitted states) to bound Eve’s information. This allows us to bound the secret key rate. Recall that here, the transmitted states can be expressed as per Equation 1, which we use to compute the inner products.
Using this technique, we simulate the secret key rate in four cases. The results can be seen in Figure 4.
  • Assuming the experimentally measured phase profiles shown in Figure 3b. Cubic spline interpolations (generated using the SciPy interpolation class) were used to represent the profiles, as shown in Figure 3b Contributors (2022). This allows the Gram matrix to be computed more reliably.
  • Assuming the optimistically simulated phase profile from Figure 1.
  • Assuming square phase profiles with a phase corresponding to the average phase of the experimentally measured phase profiles. We define the average phase ($\gamma_{avg}$) of an experimentally measured phase profiles ($\gamma(t)$) as follows.
[…]
    Recall that $f(t)$ represents the Gaussian optical intensity profile of the transmitted states.
  • Assuming ideal square BB84 profiles.
Discussion.— In Figure 4, we see that the experimentally collected phase profiles lead to a significant (over 50%) reduction in the maximum secure transmission distance, relative to the ideal case. Since the simulated phase profiles are optimistic and do not consider any imperfections beyond bandwidth limitations, they perform slightly better. However, even in this relatively optimistic case, the maximum secure transmission distance is reduced by over 50%.
Comparatively, despite having the same average encoding error as the experimentally collected phase profiles, the square profiles only cause a marginal reduction in the maximum secure transmission distance. Here comes a key conceptual point. A square phase profile implies that transmitted states satisfy the qubit assumption and hence do not carry a side channel. This results in a higher maximum secure transmission distance, as the protocol is now loss-tolerant Currás-Lorenzo et al. (2023).
In contrast, a temporally varying phase profile violates the qubit assumption. This provides additional dimensions to the Hilbert space and may enable an eavesdropper to perform new attack strategies. For this reason, the secure key is substantially reduced.
Figure 4: Key rate results (with respect to the distance between Alice/Bob and the central relay) assuming an optical intensity profile with a 50 ps width (FWHM). In blue is the key rate plot which comes from using the experimentally collected phase profiles (from Figure 3). In orange is the key rate plot which comes from using the simulated phase profile (from Figure 1). In green is the key rate plot which comes from assuming square phase profiles with a time-independent error. This error corresponds to the average error of the experimental profiles. In red is the key rate plot which comes from assuming ideal (square) phase profiles.
To further understand our key rate results, we can compute the corresponding $\epsilon^{i}_{x}$ values for each case (see Table 1). We can see that the experimentally collected and simulated phase profiles lead to $\epsilon^{i}_{x}$ values on the order of $10^{-3}$. Comparatively, the square phase profiles lead to $\epsilon^{i}_{x}$ values of 0, as they correspond to qubit states. Referring back to Currás-Lorenzo et al. (2023), we can qualitatively understand the impact of $\epsilon^{i}_{x}$ of various orders of magnitude on the secret key rate. For a BB84 protocol using single photons, Currás-Lorenzo et al. (2023) reports a reduction in the maximum secure transmission distance of over 70% (relative to the qubit case) for an $\epsilon^{i}_{x}$ on the order of $10^{-3}$. In addition to our key rate results, this further emphasizes the substantial impact which this side channel has on security.
Table 1: Table of $\epsilon^{i}_{x}$ values corresponding to the experimental and simulated cases in Figure 4.
$\pi$ $\pi/2$
exp profile $2.661\times 10^{-3}$ $8.193\times 10^{-4}$
sim profile $1.019\times 10^{-3}$ $2.556\times 10^{-4}$
Naturally, we would expect the time-variance of the phase profiles to have less of an impact on the secret key rate for smaller optical intensity pulse widths. The smaller the optical intensity pulse width, the more constant the temporal phase profile would be over the duration of the pulse width. To confirm this, we computed the secret key rate assuming optical pulse widths (FWHM) of 25 ps and 12.5 ps. The results can be seen in Figure 5. The corresponding $\epsilon^{i}_{x}$ values are also provided in Table 2. As expected, we see dramatic improvements in the secret key rate as the optical pulse width gets smaller in relation to the 200 ps modulation pulse width (along with order of magnitude changes in $\epsilon^{i}_{x}$). This suggests that this side channel can be mitigated by either shortening the optical pulse width or increasing the modulation pulse width. Increasing the modulation pulse width would shorten the duty cycle of the modulation pulse train - this would exacerbate the issue of pulse correlations, which are a major security concern in high clock rate QKD systems Pereira et al. (2020). Shortening the optical pulse width is a relatively viable option - however, for substantial improvements to the secret key rate, techniques such as mode-locking are required for generating pulses on the order of ps Martinez and Yamashita (2011); Zhao et al. (2021).
Figure 5: Key rate results (with respect to the distance between Alice/Bob and the central relay) using the experimentally collected phase profiles (from Figure 3). The results from using the simulated phase profile (from Figure 1) are shown using a dotted line. Here, the width (FWHM) of the Gaussian optical intensity profile is varied. Blue corresponds to a 50 ps width (illustrated in green in Figure 3). Orange corresponds to a 25 ps width and green corresponds a 12.5 ps width. Red corresponds to the case of ideal, time-invariant encoding.
Table 2: Table of $\epsilon^{i}_{x}$ values corresponding to the cases in Figure 5 (varying optical pulse widths).
$\pi$ $\pi/2$
50 ps $2.661\times 10^{-3}$ $8.193\times 10^{-4}$
25 ps $8.221\times 10^{-4}$ $1.684\times 10^{-4}$
12.5 ps $1.141\times 10^{-4}$ $4.120\times 10^{-5}$
Conclusion.— In this work, we have identified a side channel which arises in GHz clock rate QKD sources which make use of bandwidth-limited modulation for encoding. At such high clock rates, bandwidth limitations of source equipment start to present several security issues. One such issue is encoding which varies throughout the duration of the optical pulse. This temporal degree of freedom causes the transmitted states to break the qubit assumption, and hence acts as a source of information leakage. This is a side channel which has not been studied in the literature.
We conducted an investigation into this newly identified side channel, using BB84 MDI QKD with decoy states as an illustrative example. Using experimental and simulated results, we demonstrate that this side channel can bring down the secret key rate by several orders of magnitude, as shown in Figure 4.
Although we identified a potential mitigation method, the ideal option would be to remove the time-variance of the phase profiles entirely. At GHz repetition rates, this would be practically unfeasible when using active modulation due to bandwidth limitations. Hence, this work provides further motivation for the use of passive QKD sources, as they circumvent the need for active modulation components Wang et al. (2023); Zapatero et al. (2023); Hu et al. (2023).
Crucially, our work further demonstrates how seemingly minor practical factors in QKD implementations can lead to substantial security risks, particularly when using active modulation components. This highlights the need for experimentalists to be highly aware of imperfections in QKD experiments.
Acknowledgements.
We would like to thank Reem Mandil, Alexander Greenwood, and Andi Shahaj for their support. This research was supported in part by NSERC, CFI, ORF, MITACS, and DRDC.
Supplemental Material - Key Rate Analysis Technique
Here, we provide a high level picture for the key rate simulation technique S_Primaatmaja2019 which we employ. This technique bounds the secret key rate by bounding the single photon phase error rate ($e_{ph}$) of the QKD protocol using numerical optimization. The optimization problem can be written as follows.
maximize $e_{ph}$
s.t. $\braket{\psi^{i^{\prime}}_{x^{\prime}}\varphi^{j^{\prime}}_{y^{\prime}}}{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}=\sum_{z}\braket{e^{i^{\prime},j^{\prime}}_{x^{\prime},y^{\prime},z}}{e^{i,j}_{x,y,z}}_{E}$ (a)
$p_{\texttt{pass},L}^{i,j,x,y}\leq\braket{e^{i,j}_{x,y,P}}{e^{i,j}_{x,y,P}}_{E}\leq p_{\texttt{pass},U}^{i,j,x,y}$ (b)
$0\leq e_{ph}\leq\frac{1}{2}$ (c)
$G_{E}\succeq 0$ (d)
Here,
  • |$\ket{\psi^{i}_{x}}_{A}$ and |$\ket{\phi^{j}_{y}}_{B}$ represent the single photon component of Alice and Bob’s transmitted coherent states, respectively. Superscripts indicate the basis choice (0 denoting the key generation basis and 1 denoting the conjugate basis) while subscripts indicate the encoded bit (0 or 1). Hence, $\braket{\psi^{i^{\prime}}_{x^{\prime}}\varphi^{j^{\prime}}_{y^{\prime}}}{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}$ represent the elements of Alice and Bob’s Gramian matrix S_Primaatmaja2019 . It is through these inner products that the temporal phase profiles enter the key rate analysis.
  • |$\ket{e_{x,y,z}^{i,j}}_{E}$ (Eve’s state) can be defined as per Equation S1, through the unitary evolution of Alice and Bob’s transmitted state. Here, $z\in\{P,F\}$ represents Eve’s announcement (pass/fail) in the protocol and $\braket{z}{z^{\prime}}_{Z}=\delta_{z,z^{\prime}}$. Note that |$\ket{e^{i,j}_{x,y,z}}_{E}$ are subnormalized such that |$\ket{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}$ are normalized.
    |$\ket{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}\rightarrow\sum_{z=P,F}\ket{e^{i,j}_{x,y,z}}_{E}\ket{z}_{Z}$ (S1)
  • From Equation S1, it is apparent that $\braket{e^{i,j}_{x,y,P}}{e^{i,j}_{x,y,P}}_{E}$ refers to the passing probability of the transmitted state |$\ket{\psi^{i}_{x}\varphi^{j}_{y}}_{AB}$, since |$\ket{e^{i,j}_{x,y,z}}_{E}$ are subnormalized. The variables $p_{\texttt{pass},L}^{i,j,x,y}$ and $p_{\texttt{pass},U}^{i,j,x,y}$ refer to the upper and lower bound on this passing probability. They can be found through the passing probability of Alice and Bob’s overall transmitted coherent state S_Bourassa2022 .
  • $e_{ph}$ can be expressed as follows using the elements of Eve’s Gramian matrix S_Bourassa2022 .
    $\frac{1}{2}-\frac{Re\left(\braket{e^{0,0}_{0,0,P}}{e^{0,0}_{1,1,P}}_{E}+\braket{e^{0,0}_{0,1,P}}{e^{0,0}_{1,0,P}}_{E}\right)}{\sum_{x,y}p_{\texttt{pass},L}^{0,0,x,y}}$ (S2)
  • $G_{E}$ refers to Eve’s Gramian matrix (which has elements $\braket{e^{i^{\prime},j^{\prime}}_{x^{\prime},y^{\prime},z^{\prime}}}{e^{i,j}_{x,y,z}}_{E}$).
Overall, this is an optimization of $e_{ph}$ with respect to Eve’s Gramian matrix. The known values in this optimization problem are Alice and Bob’s transmitted states. The unknown values are $e_{ph}$ (the objective function) and Eve’s Gramian matrix, $G_{E}$.
After bounding $e_{ph}$, we can evaluate the secret key rate, $R$. The formula for the asymptotic secret key rate in decoy state MDI QKD is as follows (neglecting the effects of sifting and assuming perfect error correction efficiency) S_Ma2012_mdi_decoy .
$R\geq p_{\texttt{pass},L}^{0,0}[1-h_{2}(e_{ph})]-Qh_{2}(E_{bit}),$ (S3)
Here,
  • $p_{\texttt{pass},L}^{0,0}=\sum_{x,y}p_{\texttt{pass},L}^{0,0,x,y}$ refers to the total passing probability (lower bound) of the transmitted states (single photon component) in the key generation basis ($i,j=0$) over all possible $x,y$.
  • $Q$ refers to the passing probability of signal states in the key generation basis.
  • $E_{bit}$ is the bit error rate. Concretely, it’s the proportion of passed signal states in the key generation basis in which Alice and Bob encode opposite bits.
  • $h_{2}$ refers to the binary entropy function.